# Generalization in Reinforcement Learning for Radio Access Networks

### v2

@@ Abstract
nea-real-time RAN intelligent controller
@@ 1 INTRODUCTION
However, several studies Farebrother et al. (2018); Zhang et al. (2018b, a); Gamrian and Goldberg (2019); Song et al. (2019); Igl et al. (2019) have highlighted the limited generalization ability of RL agents to transfer learned knowledge to novel environments. This limitation primarily arises from training algorithms within fixed environments, which results in overfitting to specific transition dynamics and reward structures Farebrother et al. (2018). Consequently, such models perform well under in-distribution conditions but suffer significant degradation when exposed to unseen environments. The challenge of RL generalization is exacerbated in partially observable or uncertain Markov decision processes, where incomplete knowledge or environmental uncertainty renders available states insufficient for optimal decision-making Igl et al. (2019).
To generalize effectively, RL agents must learn transferable representations that are robust to variations in the environment, including uncertainty and feature variability Kirk et al. (2023). While classical regularization techniques from supervised learning—such as weight decay, dropout, and batch or layer normalization—can enhance the robustness of RL policies Cobbe et al. (2019), domain randomization has recently emerged as a particularly effective alternative Tobin et al. (2017). This approach systematically introduces variability into training environments—through perturbations in state parameters or environmental dynamics—encouraging agents to learn robust policies that remain effective under such variations.
This technique is particularly valuable for sim-to-real (sim2real) transfer, where it helps RL policies maintain performance when deployed in real-world environments, as demonstrated in robotics Zhao et al. (2020). In complex and dynamic domains such as RANs, where live system training is often impractical or prohibitively costly, domain randomization facilitates robust sim2real transfer. A complementary approach is sim-to-sim (sim2sim) skill transfer Gordon et al. (2019), which enables RL agents to transfer policies across simulated environments with varying characteristics, thereby enhancing generalization and increasing the likelihood of successful real-world deployment without repeated retraining Wang et al. (2021).
@@
To address this challenge, this paper investigates generalization in RL for RAN applications, with a focus on improving adaptability and robustness under heterogeneous network conditions. The key contributions of this work are as follows:
@@
  • We introduce a distributed learning architecture designed to address the dual challenge of ensuring data diversity and managing the computational complexity of training generalizable models. We show how this architecture supports large-scale data collection across simulated environments (for sim2sim transfer) and how its principles can extend to hierarchical, decentralized real-world RANs.
@@
The paper is organized as follows. Sections 2 and 3 discuss the challenges and key enablers of RL generalization in RAN applications. Section 4 introduces graph models to enrich state representations with node attributes and topological information, while Section 5 describes a distributed learning architecture designed to enhance RL generalization in both simulated and live RANs. Section 6 details the link adaptation case study, and Section 7 outlines the corresponding RL design. Finally, Section 8 presents numerical evaluations, and Section 9 concludes the paper.
@@ 2.1 Definitions and Motivations
Following LeCun et al. (2015), we define RL generalization in the RAN context as the ability of a policy to perform well under new, previously unseen network conditions, deployments, and radio environments. We focus in particular on zero-shot generalization (ZSG) Kirk et al. (2023), which aims to produce RL policies that perform well in unseen RAN environments without additional training. This capability is critical because retraining for every environmental change (e.g., traffic pattern or network load variation) or deployment difference (e.g., different cells) is costly and impractical.
@@
Achieving ZSG is also essential for scalable deployment of AI models in RAN systems. It enables network-wide replacement of a given function with a single, generalized policy, ensuring consistent performance and reducing the need for continual retraining. In contrast, relying on cell-specific fine-tuned models would degrade robustness and severely hinder scalability due to the increased complexity of lifecycle and data management. For example, maintaining separate models for thousands of radio cells for each AI-driven RAN function would be operationally burdensome and risk inconsistent behavior across the network.
@@ 2.2 Challenges
    Heterogeneity of RAN environments: RAN deployments vary significantly in topology, scale, and radio access technologies. Differences in cell types (e.g., macro vs. small cells), hardware and deployment configurations (e.g., site location, antenna type and geometry), radio technologies (e.g., MIMO, mMIMO), and user distributions contribute to a highly diverse operational landscape. Such variability poses major obstacles to RL generalization in RANs.
@@
    Non-stationary network dynamics: Wireless environments are inherently non-stationary. Factors such as user mobility, varying traffic loads, and time- and frequency-dependent channel conditions continually alter the environment’s transition and reward dynamics. Consequently, policies trained under specific traffic or mobility scenarios may degrade when deployed under sudden congestion, sparse demand, or changing user behavior. This temporal variability challenges static RL models and necessitates adaptive policy mechanisms.
@@ 3 ENABLERS OF RL GENERALIZATION IN RAN
To promote RL generalization for RAN control functions, we identify three function-agnostic enablers: (i) robust state reconstruction (to handle partial and noisy observations), (ii) training-environment diversity (to cover a broad range of deployment scenarios), and (iii) distributed learning (to scale up data collection across many environments). In subsequent sections, we demonstrate their application to RL-based link adaptation.
@@ 3.1 Robust State Reconstruction
A prerequisite to attain model generalization over the RAN environment is robustness to uncertainty, that is, the model ability to cope with epistemic uncertainty arising from incomplete knowledge or limited data regarding the training domain Igl et al. (2019). Epistemic uncertainty, for example, occurs when certain areas of the input space are poorly represented or completely missing in the training data, preventing the model from making confident predictions. To ensure robustness under uncertainty for RL applications in RAN, the input state of a model must be carefully constructed to accurately reflect environmental changes, even when the available observations are limited, delayed, or noisy.
RAN functionalities typically operate under partial observability and heterogeneous information aging. Access nodes, for example, can only rely on local or incomplete views of the network state—either from direct sensing or from measurements reported by users or neighboring nodes. Furthermore, such information is inherently affected by uncertainties in its availability when a model needs it. For instance, while a model design may require certain measurements, such data may be temporarily unavailable, outdated, or unsupported depending on the capability of specific network nodes or user devices. Moreover, different observations needed to reconstruct the state of a RAN functionality may be gathered with very diverse timescales. As a result, the state reconstructed at a given time often contains features with different aging rates, leading to incomplete representations of the true state of the underlying system.
@@
This classification allows us to consider different approaches to reconstruct the state of a RAN function. The classic approach is to feed all information, indistinctly, to a deep Q-network (DQN) model. However, to improve generalization across diverse RAN deployments and propagation environments, static and semi-static information can be processed using graph-based methods, such as graph neural networks Scarselli et al. (2009). Graph structures allow us to embed both radio access nodes and cell attributes, as well as their topological relationships, into rich state representations that can be transferred across different network deployments, improving model robustness to variations in network layout and environmental conditions Suárez-Varela et al. (2023); Shen et al. (2023); Lu et al. (2025). Section 4 further expands on this aspect.
@@ 3.2 Training Environment Diversification
An effective approach to achieving ZSG in RL is domain randomization Zhao et al. (2020), which deliberately varies aspects of the training environment so that agents learn robust, transferable policies. By exposing the agent to conditions often beyond those expected in deployment—through randomized user locations, traffic loads, channel conditions, and interference levels—the learned policy overfits less and generalizes better. In this way, domain randomization narrows the sim2real gap by making simulations sufficiently diverse to capture real-world uncertainty.
In the context of RANs, domain randomization entails sampling heterogeneous network topologies, traffic and load scenarios, mobility patterns, and user densities. For example, an RL–based scheduler could train on randomized combinations of traffic demands and cell loads to improve generalization to unpredictable deployment conditions. Prior work in robotics has demonstrated the benefits of zero-shot transfer Zhao et al. (2020), and we argue similar benefits for RL in wireless networks, particularly when combined with domain adaptation techniques.
@@
    Unrealistic extremes: Unconstrained randomization can yield non-physical scenarios—such as channel models with zero path loss or deployments involving hundreds of user equipments in a 100 meters cell—that misguide learning. To prevent this, domain knowledge and expert-defined constraints should be applied to keep randomization within physically plausible limits.
@@
    Training instability: Domain randomization can introduce high variance in reward signals, destabilizing training and leading to non-convergent policies, excessive exploration, or catastrophic forgetting. This instability can be mitigated through curriculum learning Graves et al. (2017); Matiisen et al. (2020), which gradually increases task complexity or environment randomness to ensure stable and progressive training.
@@ 3.3 Distributed Learning
Distributed learning improves model robustness in sim2sim transfer by enabling parallel exploration of diverse environmental configurations. Multiple actors interact with distinct randomized environments concurrently, achieving broader domain coverage and reducing wall-clock time compared to sequential sampling. Architectures such as IMPALA Espeholt et al. (2018), A3C Mnih et al. (2016b), Ape-X Horgan et al. (2018), and distributed PPO OpenAI et al. (2019) asynchronously update a shared policy from heterogeneous experiences, stabilizing training and mitigating overfitting to narrowly sampled conditions. A centralized replay buffer or gradient aggregator further integrates experiences across configurations, exposing the learner to a wider range of gradients and fostering robust, generalizable policies.
In Section 5, we present an architecture that scales distributed learning in RAN simulators to exploit environment diversification. We also discuss how its design principles can apply to live RAN deployments, where data diversity arises from the collective experience of distributed network nodes performing policy evaluation during training.
@@ 4 GRAPHS
GNN extend traditional neural networks to graph-structured data, allowing the incorporation of connectivity information into the learning process. To achieve this, GNNs typically employ a feedforward architecture with multiple layers that iteratively aggregate and transform information characterizing the local neighborhood of a node using learnable permutation-invariant functions Scarselli et al. (2009). At each layer of a GNN, a node updates its embedding by combining its own features with a summary of its neighbors’ features, typically through operations such as message passing and neighborhood pooling Scarselli et al. (2009). The number of layers determines how far information propagates in the graph: for instance, one layer captures immediate neighbors, two layers include two-hop neighbors, and so on.
@@
    GCN: Extend convolutional networks to graph structures, effectively capturing local node interactions Kipf and Welling (2017).
@@
    GAT: Incorporate attention mechanisms to weigh the importance of different neighbors during message passing Veličković et al. (2018).
@@
  • Graph transformers: Leverage the self-attention mechanism of transformers to model long-range dependencies in graph-structured data, potentially capturing complex hierarchical relationships in RANs systems Ying et al. (2021).
@@
Graph representations, such as those learned by GNNs, are particularly well suited for modeling RAN deployments, as they can capture both fine-grained local interactions (e.g., interference between nearby cells) and high-level global patterns (e.g., backbone connectivity or routing hierarchies) Lu et al. (2025). We adopt this approach to learn expressive network embeddings from static and semi-static network information that captures both site/cell attributes and inter-site/cell relationships within a RAN topology. Since GNN-based models share aggregation functions across all nodes and edges, the same model can be reused—without retraining—for graphs with varying topologies that represent spatial and relational patterns of different portions of a RAN deployment. Therefore, graph models facilitate better RL generalization in RAN by improving the robustness and transferability of the model across diverse deployment scenarios. Furthermore, because the information characterizing a RAN deployment is common across different functionalities, a learned embedding of the deployment can be reused in a variety of applications, including interference management, mobility management, and traffic prediction Shen et al. (2023).
@@
Different graph representations of a RAN deployment can be constructed by modeling base stations, antennas, or user devices as nodes, while edges can capture communication links, interference relationships and geographical proximity. The graph structure in GNN-based models is often tailored to the specific use case. For example, Figure 1 illustrates a graph representing a UE-centric view of a RAN neighborhood, where the central node represents the serving cell of a UE, and surrounding nodes represent relevant interfering cells. Each node $i$ is associated with a feature vector $\mathbf{h}_{i}$ that characterizes the corresponding radio cell, while edges define proximity relationships between the associated cells.
@@
The approach shown in Figure 1 is useful for designing RAN functionalities that operate on UE-specific parameters, such as link power control, transmission beamforming, or link adaptation, where the functional state varies per UE. In this case, the graph construction ensures that the context information provided to the GNN reflects the portion of the network topology most relevant to each UE—e.g., capturing relationships between surrounding cells that affect interference patterns, signal strength, and traffic conditions. In such a scenario, the serving cell can derive the graph instance using UE-side measurements (e.g., Reference Signal Receive Power (RSRP) values of neighboring cells), while additional context information can be exchanged via standardized interfaces, such as the Xn interface in 5G NR networks.
@@
Section 8 describes how we integrate this approach with a DQN design for study case presented in Section 6.
@@ 5 DISTRIBUTED LEARNING
Since the introduction of the seminal DQN algorithm Mnih et al. (2013), several algorithmic and architectural advancements—such as GORILA Nair et al. (2015), A3C Mnih et al. (2016a), IMPALA Espeholt et al. (2018), Ape-X Horgan et al. (2018), R2D2 Kapturowski et al. (2018), Seed RL Espeholt et al. (2020), NGU Badia et al. (2020b), and Agent57 Badia et al. (2020a)—have progressively improved the scalability and effectiveness of distributed RL algorithms. Notably, IMPALA demonstrated how off-policy learning, combined with distributed actor-learner architectures, can effectively mitigate latency between action generation and gradient updates. These methods employ asynchronous updates, with distributed actors concurrently generating and evaluating training data across multiple parallel environments. This approach leads to more frequent policy updates, thereby accelerating learning and significantly improving performance. For example, the R2D2 architecture achieved up to a 20-fold increase in performance and a 50-fold reduction in training time compared to traditional DQN implementations.

### v3

@@ Abstract
real-time RAN intelligent controller
near-RT RIC
near-real-time RAN intelligent controller
@@ 1 Introduction
However, several studies FMB:18; ZVM+:18; ZBP:18; GaG:19; SJT+:19; ICL+:19 have highlighted the limited generalization ability of RL agents to transfer learned knowledge to novel environments. This limitation primarily arises from training algorithms within fixed environments, which results in overfitting to specific transition dynamics and reward structures FMB:18. Consequently, such models perform well under in-distribution conditions but suffer significant degradation when exposed to unseen environments. The challenge of RL generalization is exacerbated in partially observable or uncertain Markov decision processes, where incomplete knowledge or environmental uncertainty render the available states insufficient for optimal decision-making ICL+:19.
To generalize effectively, RL agents must learn transferable representations that are robust to variations in the environment, including uncertainty and feature variability KZG+:23. While classical regularization techniques from supervised learning—such as weight decay, dropout, and batch or layer normalization—can enhance the robustness of RL policies CKH+:19, domain randomization has recently emerged as a particularly effective alternative TFR+:17. This approach systematically introduces variability into training environments—through perturbations in state parameters or environmental dynamics—encouraging agents to learn robust policies that remain effective under such variations.
This technique is particularly valuable for sim-to-real (sim2real) transfer, where it helps RL policies maintain performance when deployed in real-world environments, as demonstrated in robotics ZWQ+:20. In complex and dynamic domains such as RANs, where live system training is often impractical or prohibitively costly, domain randomization facilitates robust sim2real transfer. A complementary approach is sim-to-sim (sim2sim) skill transfer GKP+:19, which enables RL agents to transfer policies across simulated environments with varying characteristics, thereby enhancing generalization and increasing the likelihood of successful real-world deployment without repeated retraining WAB:21.
@@
To address this challenge, this paper investigates generalization in RL for RAN applications, with a focus on improving adaptability and robustness under heterogeneous network conditions. The main contributions are listed as:
@@
  • We introduce a distributed learning architecture designed to address the dual challenge of ensuring data diversity and managing the computational complexity of training generalizable models. We show how this architecture supports large-scale data collection across simulated environments (for sim2sim transfer) and how its principles can extend to hierarchical, decentralized real-world RANs to learn policies from field data.
@@
The paper is organized as follows. Sections˜2 and 3 discuss the challenges and key enablers of RL generalization in RAN applications. Section˜4 introduces graph models to enrich state representations with node attributes and topological information, while Sections˜5 and 6 describe a distributed learning architecture designed to enhance RL generalization in simulated and real-world RANs, respectively. Section˜7 details the link adaptation case study, and Section˜8 outlines the corresponding RL design. Finally, Section˜9 presents numerical evaluations, and Section˜10 concludes the paper.
@@ 2.1 Definitions and Motivations
Following DeepLearning:15, we define RL generalization in the RAN context as the ability of a policy to perform well under previously unseen network conditions, deployments, and radio environments. We focus in particular on zero-shot generalization (ZSG) KZG+:23, which aims to produce RL policies that perform well in unseen RAN environments without additional training. This capability is critical because retraining for every environmental change (e.g., traffic pattern or network load variation) or deployment difference (e.g., different cells) is costly and impractical.
@@
Achieving ZSG is also essential for scalable deployment of AI models in RAN systems. It enables network-wide replacement of a given RAN function with a single, generalized policy, ensuring consistent performance and reducing the need for continual retraining. In contrast, relying on cell-specific fine-tuned models would degrade robustness and severely hinder scalability due to the increased complexity of model lifecycle and data management. For example, maintaining separate models for thousands of radio cells for each AI-driven RAN function would be operationally burdensome and risk inconsistent behavior across the network.
@@ 2.2 Challenges
    Heterogeneity of RAN environments: RAN deployments vary significantly in topology, scale, and radio access technologies. Differences in cell types (e.g., macro vs. small cells), hardware and deployment configurations (e.g., site location, antenna type and geometry), radio technologies (e.g., MIMO, mMIMO), as well as diversity of user devices and their distributions contribute to a highly diverse operational landscape. This heterogeneity poses major obstacles to generalization.
@@
    Non-stationary network dynamics: Wireless environments are inherently non-stationary. Factors such as user mobility, varying traffic loads, and time- and frequency-dependent channel conditions continually alter the environment’s transition and reward dynamics. Consequently, policies trained under specific traffic or mobility scenarios may degrade when deployed under sudden congestion, sparse demand, or changing user behavior. This temporal variability challenges static RL models and requires adaptive mechanisms.
@@ 3 Enablers of RL Generalization in RAN
To promote RL generalization for RAN control functions, we identify three function-agnostic enablers: (i) robust state reconstruction (to handle partial and noisy observations), (ii) training-environment diversity (to cover a broad range of deployment scenarios), and (iii) distributed learning (to scale up data generation across many environments). In subsequent sections, we demonstrate their application to link adaptation.
@@ 3.1 Robust State Reconstruction
A prerequisite to attain model generalization over the RAN environment is robustness to uncertainty, that is, the model’s ability to cope with epistemic uncertainty arising from incomplete knowledge or limited data regarding the training domain ICL+:19. Epistemic uncertainty, for example, occurs when certain areas of the input space are poorly represented or completely missing in the training data, preventing the model from making confident predictions. To ensure robustness under uncertainty for RL applications in RAN, the input state of a model must be carefully constructed to accurately reflect environmental changes, even when the available observations are limited, delayed, or noisy.
RAN functionalities typically operate under partial observability and heterogeneous information aging. Access nodes, for example, can only rely on local or incomplete views of the network state—either from direct sensing or from measurements reported by users or neighboring nodes. Furthermore, such information is inherently affected by uncertainties in its availability when a model needs it. For instance, while a model design may require certain measurements, such data may be temporarily unavailable, outdated, or unsupported, depending on the capability of specific network nodes or user devices. Moreover, different observations needed to reconstruct the state of a RAN functionality may be gathered with very diverse timescales. As a result, the state reconstructed at a given time often contains features with different aging rates, leading to incomplete representations of the true state of the underlying system.
@@
This classification allows us to consider different approaches to reconstruct the state of a RAN function. The classic approach is to feed all information, indistinctly, to a deep Q-network (DQN) model. However, to improve generalization across diverse RAN deployments and propagation environments, static and semi-static information can be processed using graph-based methods, such as graph neural networks SGT:09. Graph structures allow us to embed both radio access nodes and cell attributes, as well as their topological relationships, into rich state representations that can be transferred across different network deployments, improving model robustness to variations in network layout and environmental conditions SAF:23; SZS:23; LLZ:25. Section˜4 further expands on this aspect.
@@ 3.2 Training Environment Diversification
An effective approach to achieving ZSG in RL is domain randomization ZWQ+:20, which deliberately varies aspects of the training environment so that agents learn robust, transferable policies. By exposing the agent to conditions often beyond those expected in deployment—through randomized user locations, traffic loads, channel conditions, and interference levels—the learned policy overfits less and generalizes better. In this way, domain randomization narrows the sim2real gap by making simulations sufficiently diverse to capture real-world uncertainty.
In the context of RANs, domain randomization entails sampling heterogeneous network topologies, traffic and load scenarios, mobility patterns, and user densities. For example, an RL–based scheduler could train on randomized combinations of traffic demands and cell loads to improve generalization to unpredictable deployment conditions. Prior work in robotics has demonstrated the benefits of zero-shot transfer ZWQ+:20, and we argue similar benefits for RL in wireless networks, particularly when combined with domain adaptation techniques.
@@
    Unrealistic extremes: Unconstrained randomization can yield non-physical scenarios—such as channel models with zero path loss or deployments involving hundreds of user equipments within a $100$-$\mathrm{m}\text{/}$ cell—that misguide learning. To prevent this, domain knowledge and expert-defined constraints should be applied to keep randomization within physically plausible limits.
@@
    Training instability: Domain randomization can introduce high variance in reward signals, destabilizing training and leading to non-convergent policies, excessive exploration, or catastrophic forgetting. This instability can be mitigated through curriculum learning GBM+:17; MOC:20, which gradually increases task complexity or environment randomness to ensure stable and progressive training.
@@ 3.3 Distributed Learning
Distributed learning improves model robustness in sim2sim transfer by enabling parallel exploration of diverse environmental configurations. Multiple actors interact with distinct randomized environments concurrently, achieving broader domain coverage and reducing wall-clock time compared to sequential sampling. Architectures such as IMPALA Espeholt:2018, A3C MBM+:16, Ape-X Horgan:2018, and distributed PPO openai2019dota2largescale asynchronously update a shared policy from heterogeneous experiences, stabilizing training and mitigating overfitting to narrowly sampled conditions. A centralized replay buffer or gradient aggregator further integrates experiences across configurations, exposing the learner to a wider range of gradients and fostering robust, generalizable policies.
Section˜5 presents an architecture that scales distributed learning in RAN simulators to leverage environmental diversity. Section˜6 extends these design principles to real-world RAN deployments, where data diversity stems from the collective experience of distributed network nodes performing policy evaluation concurrently during training.
@@ 4 Graphs
GNN extend traditional neural networks to graph-structured data, allowing the incorporation of connectivity information into the learning process. To achieve this, GNNs typically employ a feedforward architecture with multiple layers that iteratively aggregate and transform information characterizing the local neighborhood of a node using learnable permutation-invariant functions SGT:09. At each layer of a GNN, a node updates its embedding by combining its own features with a summary of its neighbors’ features, typically through operations such as message passing and neighborhood pooling SGT:09. The number of layers determines how far information propagates in the graph: for instance, one layer captures immediate neighbors, two layers include two-hop neighbors, and so on.
@@
    GCN: Extend convolutional networks to graph structures, effectively capturing local node interactions KW:17.
@@
    GAT: Incorporate attention mechanisms to weigh the importance of different neighbors during message passing VCC:18.
@@
  • Graph transformers: Leverage the self-attention mechanism of transformers to model long-range dependencies in graph-structured data, potentially capturing complex hierarchical relationships in RANs systems YCL:21.
Graph representations, such as those learned by GNNs, are particularly well suited for modeling RAN deployments, as they can capture both fine-grained local interactions (e.g., interference between nearby cells) and high-level global patterns (e.g., backbone connectivity or routing hierarchies) LLZ:25.
@@
We adopt this approach to learn expressive network embeddings from static and semi-static network information that captures both site/cell attributes and inter-site/cell relationships within a RAN topology. Since GNN-based models share aggregation functions across all nodes and edges, the same model can be reused—without retraining—for graphs with varying topologies that represent spatial and relational patterns of different portions of a RAN deployment. Therefore, graph models can facilitate better RL generalization in RAN by improving the robustness and transferability of the model across diverse deployment scenarios. Furthermore, because the information characterizing a RAN deployment is common across different functionalities, a learned embedding of the deployment can be reused in a variety of applications, including interference management, mobility management, and traffic prediction SZS:23.
@@
Different graph representations of a RAN deployment can be constructed by modeling base stations, antennas, or user devices as nodes, while edges can capture communication links, interference relationships, and geographical proximity. The graph structure in GNN-based models is often tailored to the specific use case. For example, Figure˜1 illustrates a graph representing a UE-centric view of a RAN neighborhood, where the central node represents the serving cell of a UE, and surrounding nodes represent relevant interfering cells. Each node $i$ is associated with a feature vector $\mathbf{h}_{i}$ that characterizes the corresponding radio cell, while edges define proximity relationships between the associated cells.
@@
The approach shown in Figure˜1 is useful for designing RAN functionalities that operate on UE-specific parameters, such as power control, transmission beamforming, or link adaptation, where the functional state varies per UE. In this case, the graph construction ensures that the context information provided to the GNN reflects the portion of the network topology most relevant to each UE—e.g., capturing relationships between surrounding cells that affect interference patterns, signal strength, and traffic conditions. In such a scenario, the serving cell can derive the graph instance using UE-side measurements (e.g., Reference Signal Receive Power (RSRP) values of neighboring cells), while additional context information can be exchanged via standardized interfaces, such as the Xn interface in 5G NR networks.
@@
Section˜9 describes how we integrate this approach with a DQN design for the study case presented in Section˜7.
@@ 5 Distributed Learning
Since the introduction of the seminal DQN algorithm (Mnih:2013), several algorithmic and architectural advancements—such as GORILA (Nair:2015), A3C (MBM+:16), IMPALA (Espeholt:2018), Ape-X (Horgan:2018), R2D2 (Kapturowski:2018), Seed RL (Espeholt:2020), NGU (BSV+:20), and Agent57 (BPK+:20)—have progressively improved the scalability and effectiveness of distributed RL algorithms. Notably, IMPALA demonstrated how off-policy learning, combined with distributed actor-learner architectures, can effectively mitigate latency between action generation and gradient updates. These methods employ asynchronous updates, with distributed actors concurrently generating and evaluating training data across multiple parallel environments. Such architectures accelerate convergence through more frequent policy updates, increase sample diversity by learning from multiple environments concurrently, and improve final performance, as demonstrated in Horgan:2018; Kapturowski:2018. For example, the R2D2 architecture achieved up to a 20-fold increase in performance and a 50-fold reduction in training time compared to traditional DQN implementations.